1) find the value of a for which the graph of y=ab^x is a horizontal line 2) find the value of b for which the graph of y=ab^x is a horizontal line. explain how you found your answers
Found 2 solutions by richard1234, stanbon:
Answer by richard1234(7193)

1 - if a = 0, then y = 0, which is horizontal.

2 - if b = 0, then a*b^0 = a for nonzero b. Since a is constant, the function is constant and the line is horizontal.

Answer by stanbon(75887)

1) find the value of a for which the graph of y=ab^x is a horizontal line
y = 0 when a = 0
-------------------------
2) find the value of b for which the graph of y=ab^x is a horizontal line.
y = 0 when b = 0
===================
Cheers,
Stan H.
====================

explain how you found your answers